# MaLA-500: Massive Language Adaptation of Large Language Models

### v1

@@ Abstract
Large language models have advanced the state of the art in natural language processing. However, their predominant design for English or a limited set of languages creates a substantial gap in their effectiveness for low-resource languages. To bridge this gap, we introduce MaLA-500, a novel large language model designed to cover an extensive range of 534 languages. To train MaLA-500, we employ vocabulary extension and continued pretraining on LLaMA 2 with Glot500-c. Our experiments on SIB-200 show that MaLA-500 achieves state-of-the-art in-context learning results. We release MaLA-500 at https://huggingface.co/MaLA-LM.
**footnotetext: Equal contribution.
@@ 1 Introduction
Large Language Models (LLMs), e.g., LLaMA (Touvron et al., 2023a, b), Mistral (Jiang et al., 2023, 2024), and ChatGPT, [^1] https://openai.com/blog/chatgpt have shown remarkable performance in natural language understanding and generation. Follow-up studies (Bang et al., 2023; Lai et al., 2023; Ahuja et al., 2023a, b) observe that these English-centric LLMs, such as LLaMA with mainly English as the training data, are capable of handling some high-resource non-English languages, benefiting from the inclusion of non-English language data during pretraining. However, their applicability to low-resource languages is still limited due to data scarcity and model capability in terms of model size.
@@
Recently, several generative multilingual LLMs, such as XGLM (Lin et al., 2021), mGPT (Shliazhko et al., 2022), and BLOOM (Scao et al., 2022), have emerged. Notably, the current language coverage for these LLMs is limited to up to 60 languages, highlighting the remaining need for further work on massively multilingual LLMs.
@@
ImaniGooghari et al. (2023) have achieved a significant milestone in the realm of massive language adaptation by extending the language coverage of a small-scale multilingual language model, XLM-R (Conneau et al., 2020) - an auto-encoding model with 278M parameters, from 100 languages to an impressive number of 534 languages, and introducing an extended model, Glot500-m with 395M parameters. ImaniGooghari et al. (2023) introduce the Glot500-c corpora spanning 534 languages from 47 language families, and then apply vocabulary extension and continued pretraining to create Glot500-m. The introduction of Glot500-c significantly mitigates the challenge of data scarcity for low-resource languages. Moreover, the adaptation method is more favorable than training from scratch, as it requires fewer computational resources and emits a smaller carbon footprint. This success serves as a strong motivation for our exploration into the massive language adaptation of LLMs.
@@
This work aims to extend the capabilities of LLMs to encompass a wider range of languages. Existing works on language adaptation of pretrained models provide extended coverage across a wide linguistic spectrum but are limited to relatively small model sizes - mostly at the hundred million scale. Our study pushes the boundaries by exploring language adaptation techniques for LLMs with model parameters scaling up to 10B. Our investigation delves into generative LLMs with a substantial increase in model parameters and their in-context learning capabilities that enhance contextual and linguistic relevance across a diverse range of languages. We address the challenges of adapting LLMs to low-resource languages, such as data sparsity, domain-specific vocabulary, and linguistic diversity. We study continued pretraining of open LLM (i.e., LLaMA 2 (Touvron et al., 2023b)), vocabulary extension, and adaptation techniques (i.e., LoRA low-rank reparameterization (Hu et al., 2022)). We train and release MaLA-500 that covers more than 500 languages in various domains. We evaluate MaLA-500 on SIB-200 and the results show that MaLA-500 outperforms existing open LLMs of close or slightly larger model size. This work broadens the accessibility of LLMs, making them valuable for a more diverse set of language-specific use cases, especially for low-resource ones, and addressing the equality issue by removing language barriers for speakers of many languages, especially those underrepresented languages covered by existing LLMs.
@@ 2 Massive Language Adaptation
The principle of massive language adaption of large language models accommodates the utilization of a massively multilingual corpus (Section 2.1), the strong base LLM (Section 2.2), and the technique for effective language adaption (Sections 2.3 and 2.4).
@@ 2.1 Data
We use Glot500-c (ImaniGooghari et al., 2023) covering 534 languages [^2] We define languages using the ISO 639-3 code combined with the corresponding written script. For example, “eng_Latn” represents English written in the Latin script. as the training data of MaLA-500. See §A for the list of languages with their data amounts. We sample languages from the imbalanced dataset according to a multinomial distribution, with $\alpha=0.3$ for vocabulary extension and continued pretraining.
@@ 2.2 Model
We choose LLaMA 2 (Touvron et al., 2023b) to start continual training. LLaMA series models (Touvron et al., 2023a), with model weights released publicly, have gained popularity in the research community. Despite being English-centric compared to their multilingual counterparts, they have shown remarkable capacity for multiple languages (Zhao et al., 2024). We choose the latest LLaMA 2, trained on 2 trillion tokens, as our base model to benefit from its outstanding language capacity. Our study chooses the 7B model with 32 transformer layers.
@@ 2.3 Vocabulary Extension
The original LLaMA 2’s 32,000 tokenizer covers English and a small fraction of other European languages using Latin or Cyrillic scripts. To enhance its capability and encoding efficiency for a broader range of languages, we extend the vocabulary with Glot500-c. Specifically, we initially train a multilingual tokenizer with SentencePiece (Kudo and Richardson, 2018) on the sampled Glot500-c with a vocabulary of 250,000. Subsequently, we merge the trained tokenizer with the original LLaMA 2 tokenizer by taking the union of their vocabularies. As a result, we obtain the MaLA-500’s tokenizer with a vocabulary size of 260,164. After vocabulary extension and resizing the embedding layer, the model size becomes 8.6B.
@@ 2.4 Continued Pretraining
We employ continued pretraining for language adaptation with low-rank adaptation (LoRA, Hu et al., 2022) to enable parameter-efficient training, given the limitation of our computing resources. LoRA injects trainable rank decomposition matrices, which approximate the large weight matrices with a lower rank, to the pretrained model weights. It significantly reduces the computational complexity and thus saves the training cost. We continually train the casual language model to update the rank-decomposition matrices while freezing the model weights of pretrained models, allowing the continually trained language model to learn from new data in new languages without completely losing its previous language capacity. Continual training of large language models requires substantial computational resources. We adopt efficient distributed training setups on supercomputers to make the training process feasible.
@@ 2.5 Training
We train our model on the computing cluster with the theoretical peak performance of 2 petaflops on GPU nodes. We deploy distributed training on 24 Nvidia Ampere A100 GPUs. As for software, we utilize the Huggingface Transformers (Wolf et al., 2020), PEFT (Parameter-Efficient Fine-Tuning) [^3] https://huggingface.co/docs/peft/index, and DeepSpeed (Rasley et al., 2020). We use the ZeRO redundancy optimizer (Rajbhandari et al., 2020) and maximize the batch size that fits the memory of each GPU. We employ mixed-precision training using the bfloat16 format.
@@
The learning rate is set at 2e-4. A weight decay of 0.01 is applied to penalize large weights and mitigate overfitting. The trainable LoRA module targets all linear layers, excluding the language model head. In our setting, the final model has 10B parameters in total, in which 2B parameters are trainable. The LoRA module is incorporated with a rank of 8, an alpha value of 32, and a dropout rate of 0.05, contributing to the model’s adaptability and regularization during training. The context window is 2k. We maximize the batch size to fit the memory, making a global batch size of 384. The model undergoes three training epochs. Checkpoints are saved every 500 steps, and we employ early stopping to select the checkpoint that exhibits the most favorable average performance on downstream tasks. In this model release, we only consider the topic classification task.
@@
We train our model on CSC’s carbon-neutral data center, with all electricity generated with renewable hydropower, [^4] https://www.csc.fi/sustainable-development and the waste heat is utilized in district heating to further reduce CO2 footprint. [^5] https://www.csc.fi/sustainability-1
@@ 3 Evaluation
We evaluate the few-shot learning capability of MaLA-500 and compare it with other LLMs on a topic classification dataset of SIB-200 (Adelani et al., 2023). The classification task involves seven classes, namely science/technology, travel, politics, sports, health, entertainment, and geography. Our evaluation spans a diverse set of 177 languages, obtained by intersecting the language sets of SIB-200 and Glot500-c. In this evaluation, we compute the accuracy of 99 samples for each language, with shots randomly sampled from the respective training sets.
The evaluated LLM receives a structured prompt, which is the concatenation of few-shot examples and the sample intended for prediction. The format for both a few-shot example and the sample to predict is defined as follows:
Topic Classification: science/technology, travel, politics, sports, health, entertainment, geography.
The label of [sent] is [label]
where [sent] is the sentence for topic classification, and [label] is the ground truth. [label] is included if the sample serves as a few-shot example or omitted if it is the sample to predict. The constructed prompt is then used as input to the LLM. Subsequently, the evaluated LLM is prompted to estimate the probability of the label over the label set based on the provided prompt. This evaluation process is implemented using the lm-evaluation-harness. [^6] https://github.com/EleutherAI/lm-evaluation-harness
Table 1: 3-shot in-context learning macro-average accuracy (%) on SIB-200 of different LLMs. mGPT has no model with around 7B parameters, so we choose a larger one with 13B parameters.
Model Accuracy
LLaMA 2-7B 41.94
mGPT-13B 39.96
BLOOM-7B1 44.22
XGLM-7.5B 47.53
MaLA-500 54.10
[Uncaptioned image] Figure 1: Detailed results of 3-shot in-context learning on SIB-200. X-axis: the number of languages in different accuracy ranges (%).
Table 1: 3-shot in-context learning macro-average accuracy (%) on SIB-200 of different LLMs. mGPT has no model with around 7B parameters, so we choose a larger one with 13B parameters.
@@ 3.1 Comparison across LLMs
We compare MaLA-500 with LLaMA 2-7B, mGPT-13B, BLOOM-7B1, and XGLM-7.5B, focusing on the 3-shot in-context learning macro-average accuracy across languages. As shown in Table. 1, MaLA-500 outperforms previous LLMs significantly. Notably, MaLA-500 exhibits a remarkable 12.16% improvement compared to our base model, LLaMA 2, highlighting MaLA-500’s substantial contribution to enhancing the multilingual capacity of LLMs.
Fig. 1 provides a detailed performance analysis across languages. Notably, MaLA-500 surpasses previous LLMs in the sense that no language obtains random results (0-20%) and the number of languages with very low accuracy (20-40%) is significantly reduced. Moreover, MaLA-500 excels in the number of languages achieving 60% accuracy or higher (65 altogether).

### v2

@@ Abstract
Large language models (LLMs) have advanced the state of the art in natural language processing. However, their predominant design for English or a limited set of languages creates a substantial gap in their effectiveness for low-resource languages. To bridge this gap, we introduce MaLA-500, a novel large language model designed to cover an extensive range of 534 languages. To train MaLA-500, we employ vocabulary extension and continued pretraining on LLaMA 2 with Glot500-c. Our intrinsic evaluation demonstrates that MaLA-500 is better at predicting the given texts of low-resource languages than existing multilingual LLMs. Moreover, the extrinsic evaluation of in-context learning shows that MaLA-500 outperforms previous LLMs on SIB200 and Taxi1500 by a significant margin, i.e., 11.68% and 4.82% marco-average accuracy across languages. We release MaLA-500 at https://huggingface.co/MaLA-LM.
@@ 1 Introduction
Large Language Models (LLMs), e.g., LLaMA (Touvron et al., 2023a; b), Mistral (Jiang et al., 2023; 2024), and ChatGPT, [^1] https://openai.com/blog/chatgpt have shown remarkable performance in natural language understanding and generation. Follow-up studies (Bang et al., 2023; Lai et al., 2023; Ahuja et al., 2023a; b) observe that these English-centric LLMs, such as LLaMA with mainly English as the training data, are capable of handling some high-resource non-English languages, benefiting from the inclusion of non-English language data during pretraining. However, their applicability to low-resource languages is still limited due to data scarcity.
@@
Previous studies have released pretrained multilingual models with mostly encoder-only transformer architectures, e.g., multilingual BERT (Devlin et al., 2019) and XLM-R (Conneau et al., 2020), for around 100 languages. The paradigm shift from encoder-only to decoder-only achieves scalability for large language models with billions of model parameters, leading to the development of open multilingual models. Recently, several generative multilingual LLMs, such as XGLM (Lin et al., 2021), mGPT (Shliazhko et al., 2022), and BLOOM (Scao et al., 2022), have emerged. Notably, the current language coverage for these generative LLMs is limited to up to 60 languages, highlighting the remaining need for further work on massively multilingual LLMs for many natural languages.
@@
ImaniGooghari et al. (2023) have achieved a significant milestone in the realm of massive language adaptation by extending the language coverage of a small-scale multilingual language model, XLM-R (Conneau et al., 2020) - an auto-encoding model with 278M parameters, from 100 languages to an impressive number of 534 languages, and introducing an extended model, Glot500-m with 395M parameters. ImaniGooghari et al. (2023) introduce the Glot500-c corpora spanning 534 languages from 47 language families, and then apply vocabulary extension and continued pretraining to create Glot500-m. The introduction of Glot500-c mitigates the challenge of data scarcity for low-resource languages. Moreover, the adaptation method is more favorable than training from scratch, as it requires fewer computational resources and emits a smaller carbon footprint. This success serves as a strong motivation for our exploration into the massive language adaptation of LLMs.
@@
This work aims to extend the capabilities of LLMs to encompass a wider range of languages. Existing works like ImaniGooghari et al. (2023) on language adaptation of pretrained models provide extended coverage across a wide linguistic spectrum but are limited to relatively small model sizes - mostly at the hundred million scales, while other works like Yong et al. (2022) extended generative LLMs but are limited to a small number of languages. Our study pushes the boundaries by exploring language adaptation techniques for LLMs with model parameters scaling up to 10 billion for 534 languages. Our investigation delves into generative LLMs with a substantial increase in model parameters and their in-context learning capabilities in diverse languages, especially low-resource languages. This augmentation enables us to enhance contextual and linguistic relevance across a diverse range of languages.
We address the challenges of adapting LLMs to low-resource languages, such as data sparsity, domain-specific vocabulary, and linguistic diversity. Specifically, we study continued pretraining of open LLM, i.e., LLaMA 2 (Touvron et al., 2023b), vocabulary extension, and adaptation techniques, i.e., LoRA low-rank reparameterization (Hu et al., 2022). We deploy distributed training and release MaLA-500 that covers more than 500 languages in various domains. We evaluate MaLA-500 using intrinsic measures on held-out Glot500-c test set and parallel data and extrinsic metrics on downstream benchmarks: SIB200 and Taxi1500. The results show that MaLA-500 outperforms existing open LLMs of close or slightly larger model size. This work broadens the accessibility of LLMs, making them valuable for a more diverse set of language-specific use cases, especially for low-resource ones, and addressing the equality issue by removing language barriers for speakers of many languages, especially those underrepresented languages covered by existing LLMs.
@@ 2 Massive Language Adaptation
The principle of massive language adaptation of large language models accommodates the utilization of a massively multilingual corpus (Section 2.1), the strong base LLM (Section 2.2), and the technique for effective language adaptation: vocabulary extension (Section 2.3) and continued pretraining (Section 2.4).
@@ 2.1 Data
We use Glot500-c (ImaniGooghari et al., 2023) covering 534 languages [^2] We define languages using the ISO 639-3 code combined with the corresponding written script. For example, “eng_Latn” represents English written in the Latin script. as the training data of MaLA-500. See §A for the list of languages with their data amounts. The original number of sentences ranges from 10 thousand to 63 million. Note that Glot500-c does not put full effort into collecting data for high-resource languages but focuses on low-resource languages. We sample languages from the imbalanced dataset according to a multinomial distribution, with $\alpha=0.3$ for vocabulary extension and continued pretraining. We use different scales for sampling data to be used in model training and vocabulary construction. After sampling, the number of sentences for training ranges from 600 thousand to 8 million per language, leading to 1 billion sentences in total. The number of sentences for vocabulary construction ranges from 30 thousand to 400 thousand, making a total of 50 million sentences.
@@ 2.2 Model
We choose LLaMA 2 (Touvron et al., 2023b) to start continual training. LLaMA series models (Touvron et al., 2023a), with model weights released publicly, have gained popularity in the research community. Despite being English-centric compared to their multilingual counterparts, they have shown remarkable capacity for multiple languages (Ahuja et al., 2023b). We choose the latest LLaMA 2, trained on 2 trillion tokens, as our base model to benefit from its outstanding language capacity. Our study chooses the 7B model with 32 transformer layers, and leaves the extension of LLMs with larger sizes as a future work.
@@ 2.3 Vocabulary Extension
The original LLaMA 2’s 32,000 tokenizer covers English and a small fraction of other European languages using Latin or Cyrillic scripts. To enhance its capability and encoding efficiency for a broader range of languages, we extend the vocabulary with Glot500-c. Specifically, we initially train a multilingual tokenizer with SentencePiece (Kudo & Richardson, 2018) on the sampled Glot500-c with a vocabulary of 250,000. Subsequently, we merge the trained tokenizer with the original LLaMA 2 tokenizer by taking the union of their vocabularies. As a result, we obtain the MaLA-500’s tokenizer with a vocabulary size of 260,164. After vocabulary extension and resizing the embedding layer, the model size becomes 8.6B.
@@ 2.4 Continued Pretraining
We employ continued pretraining for language adaptation with low-rank adaptation (LoRA, Hu et al., 2022) to enable parameter-efficient training, given the limitation of our computing resources. LoRA injects trainable rank decomposition matrices, which approximate the large weight matrices with a lower rank, to the pretrained model weights. It reduces the computational complexity and thus saves the training cost while retaining high model quality (Hu et al., 2022). We continually train the casual language model to update the rank-decomposition matrices, embedding layer, and language modeling head while freezing the transformer weights of pretrained models, allowing the continually trained language model to learn from new data in new languages without completely losing its previous language capacity. Continual training of large language models requires substantial computational resources. We adopt efficient distributed training setups on supercomputers to make the training process feasible.
@@ 2.5 Training
We train our model on the computing cluster with the theoretical peak performance of 2 petaflops on GPU nodes. We deploy distributed training on 24 Nvidia Ampere A100 GPUs. As for software, we utilize the Huggingface Transformers (Wolf et al., 2020), PEFT (Parameter-Efficient Fine-Tuning), [^3] https://huggingface.co/docs/peft/index and DeepSpeed (Rasley et al., 2020). We use the ZeRO redundancy optimizer (Rajbhandari et al., 2020) and maximize the batch size that fits the memory of each GPU. We employ mixed-precision training using the bfloat16 format.
@@
The learning rate is set at 3e-4. A weight decay of 0.01 is applied to penalize large weights and mitigate overfitting. The trainable LoRA module targets the query and value matrices. The language model head is not decomposed by a LoRA module but is trained in a full-parameter manner. In our setting, the final model has 10B parameters in total, in which 2B parameters are trainable. The LoRA module is incorporated with a rank of 8, an alpha value of 32, and a dropout rate of 0.05, contributing to the model’s adaptability and regularization during training. The context window is 4k. We maximize the batch size to fit the memory, making a global batch size of 384. The model undergoes three training epochs. Checkpoints are saved every 500 steps, and we employ early stopping to select the checkpoint that exhibits the most favorable average performance on downstream tasks.
@@
We train our model on a carbon-neutral data center, with all electricity generated with renewable hydropower, and the waste heat is utilized in district heating to further reduce CO2 footprint. [^4] https://www.csc.fi/sustainable-development
@@ 3 Evaluation
3.1 Benchmarks and Setup
We consider both intrinsic and extrinsic measures for evaluation. Evaluation dataset statistics are shown in Table 1.
[…]
Intrinsic Glot500-c test (ImaniGooghari et al., 2023) $NLL$ 1000 534 Misc
PBC (Mayer & Cysouw, 2014) $NLL$ 500 370 Bible
Extrinsic SIB200 (Adelani et al., 2023) ACC 204 177 Misc
Taxi1500 (Ma et al., 2023) ACC 111 351 Bible
[…]
For intrinsic evaluation, perplexity is not comparable across models and languages due to different text segmentations. Inspired by Xue et al. (2022); Yu et al. (2023), we instead measure the negative log-likelihood ($NLL$) of the text using the given LLMs.
We concatenate the dataset as the input text and adopt the sliding-window strategy. [^5] https://huggingface.co/docs/transformers/en/perplexity The evaluation of different LLMs uses the same data with the concatenation of sentences per language, thus making $NLL$ model-comparable. In addition, we consider language-comparable $NLL$ by measuring $NLL$ on parallel data, in which every sample in different languages contains the same semantic information. We report the model-comparable $NLL$ of Glot500-c test set covering all 534 considered languages (§3.2), and language-comparable $NLL$ on Parallel Bible Corpus  (PBC, Mayer & Cysouw, 2014), covering 370 languages (§3.3).
For extrinsic evaluation, we evaluate the few-shot learning capability of MaLA-500 and compare it with other LLMs on SIB200 (Adelani et al., 2023) and Taxi1500 (Ma et al., 2023).
SIB200 is a topic classification dataset. The classification task involves seven classes, namely science/technology, travel, politics, sports, health, entertainment, and geography. Our evaluation spans a diverse set of 177 languages, obtained by intersecting the language sets of SIB200 and Glot500-c. Note that the flores200-based SIB200 evaluation set is included in the training data since Glot500-c includes flores200, but the classification labels are not provided.
Taxi1500 is another text classification dataset spanning 351 languages. It involves six classes, namely, Recommendation, Faith, Description, Sin, Grace, and Violence. Our evaluation efforts aim to cover as many languages as possible. However, the evaluation of massively multilingual language models is a challenging task. Due to the lack of real-world multilingual evaluation benchmarks, we use this benchmark that contains religious content.
For in-context learning evaluation, the evaluated LLM receives a structured prompt, which is the concatenation of few-shot examples and the sample intended for prediction. The format for both a few-shot example and the sample to predict is defined as follows:
Template for SIB200:
The topic of the news [sent] is [label]
Template for Taxi1500:
The topic of the verse [sent] is [label]
where [sent] is the sentence for classification, and [label] is the ground truth. [label] is included when the sample serves as a few-shot example but is omitted when predicting the sample. The constructed prompt is then used as input to the LLM. Subsequently, the evaluated LLM is prompted to estimate the probability of the label over the label set based on the provided prompt.
For SIB200, few-shots examples are randomly sampled from the in-language training sets. Since randomly selecting few-shot examples for in-context learning yields random results for both MaLA-500 and previous LLMs on Taxi1500, we consider the retriever-based in-context learning (Liu et al., 2022). Specifically, we use average word embeddings in layer 8 of the Glot500 (ImaniGooghari et al., 2023) for retrieving semantic-similar samples as suggested in previous work (Sabet et al., 2020) for all the compared models. The evaluation process is implemented using the lm-evaluation-harness, [^6] https://github.com/EleutherAI/lm-evaluation-harness and we use accuracy (ACC) to measure the performance of classification.
@@ 3.2 Comparison across LLMs
We compare MaLA-500 with LLaMA 2-7B, mGPT-13B, BLOOM-7B1, and XGLM-7.5B on Glot500-c test set, SIB200, Taxi1500 by computing the averaged performance across languages, and the result are given in Table 2. Among the evaluated LLMs, LLaMA 2-7B performs second-best, indicating that LLaMA 2-7B has a strong multilingual capacity and that it is reasonable to select it as the base model. MaLA-500 outperforms all compared LLMs with a close or slightly larger model size across all the evaluated tasks. Notably, compared to LLaMA 2-7B, MaLA-500 gains a lower $NLL$ on the Glot500-c test set by 39.33, and has 14.94% and 4.82% improvements on SIB200 and Taxi1500, respectively. It highlights MaLA-500’s substantial contribution to enhancing the multilingual capacity of LLMs.
Model Glot500-c test ($NLL$downarrow) SIB200 (ACC $\uparrow$) Taxi1500 (ACC $\uparrow$)
LLaMA 2-7B 190.58 42.08 44.07
mGPT-13B 282.46 45.34 40.98
BLOOM-7B1 202.95 44.63 43.98
XGLM-7.5B 205.07 34.36 43.24
MaLA-500 151.25 57.02 48.89
Table 2: Averaged results across languages on Glot500-c test (measured by $NLL$), SIB200, and Taxi1500 (measured by accuracy (%)) of different LLMs. mGPT has no model with around 7B parameters, so we choose a larger one with 13B parameters. $\downarrow$ indicates the lower, the better. $\uparrow$ indicates the higher, the better. The best results are bold.
Figures 1, 2 and 3 provide detailed performance analysis across languages on Glot500-c test, SIB200, and Taxi1500. In those figures, we group scores into different performance bins and display them in different colors. For Glot500-c test, MaLA-500 has more languages achieving better $NLL$, i.e., 61 languages with $NLL$ less than 100 and 171 languages with $NLL$ between 100 and 150. Besides, MaLA-500 has 54 (10%) languages achieving $NLL$ larger than 200, which may indicate the languages are not well covered by the measured LLM. Nevertheless, the number is much less than other LLMs. For example, the second-best LLM, LLaMA 2-7B, has 231 (43%) languages achieving $NLL$ larger than 200. For both SIB200 and Taxi1500, MaLA-500 surpasses previous LLMs in the sense that it obtains random results in fewer languages and achieves impressive performance in more languages than its counterparts.
Figure 1: $NLL$ (lower is better) on Glot500-c test with the scores grouped into four bins displayed in different colors. X-axis: the number of languages in performance ranges.
Figure 2: Accuracy (higher is better) on SIB200 with the scores grouped into four bins displayed in different colors. X-axis: the number of languages in performance ranges (%).
Figure 3: Accuracy (higher is better) on Taxi1500 with the scores grouped into four bins displayed in different colors. X-axis: the number of languages in performance ranges (%).
3.3 Comparison across Languages
To check in detail how MaLA-500 performs across languages, we check the performance across language families [^7] We assign languages to families based on Glottolog: https://glottolog.org/glottolog/family. shown in Table 3. We observe that more high-resource language families, e.g., Indo-European (indo1319) and Dravidian (drav1251), achieve slightly better performance than low-resource language families, e.g., Sino-Tibetan (sino1245).
[…]
indo1319 988M 145.35 63.53 53.03
drav1251 135M 131.29 56.25 54.65
aust1307 113M 147.37 62.83 49.69
turk1311 109M 161.71 57.08 52.55
afro1255 100M 165.46 52.00 43.74
atla1278 57M 141.92 42.90 45.52
ural1272 50M 137.52 66.67 48.58
sino1245 29M 155.64 49.30 49.31
other 60M 167.69 55.74 46.67
[…]
In Table 4, we present a comprehensive analysis of the top 5 performance improvements and declines across languages on SIB200 from MaLA-500 compared to LLaMA 2-7B. We observe that MaLA-500 has substantial improvements on low-resource scripts, e.g., Kannada (kan_Knda), while has worse performance on high-resource languages, e.g., Swedish (swe_Latn), which have been well covered by LLaMA 2-7B.
high end low end
Language LLaMA 2-7B MaLA-500 $\Delta$ Language LLaMA 2-7B MaLA-500 $\Delta$
kan_Knda 17.16 57.35 40.19 swe_Latn 71.08 60.29 -10.79
ckb_Arab 19.61 60.29 40.68 rus_Cyrl 71.57 65.20 -06.37
asm_Beng 17.16 58.82 41.66 dan_Latn 69.12 63.24 -05.88
pan_Guru 14.22 58.82 44.60 pol_Latn 74.51 68.63 -05.88
sin_Sinh 15.20 60.29 45.09 ukr_Cyrl 71.57 65.69 -05.88
Table 4: Results for five languages each with the largest (high end) and smallest (low end) gains from MaLA-500 vs. LLaMA 2-7B for SIB200. $\Delta$ indicates the difference between the scores of MaLA-500 and LLaMA 2-7B. See §B for detailed results for each task.